OpenStudy (anonymous):

An adult's ticket is two times the cost of a child's ticket. The cost for two adults' tickets and three children's tickets was $28. How much are each ticket?

OpenStudy (mertsj):

a=2c 2a+3c=28 replace a in the second equation with 2c from the first equation and solve. can you do that?

OpenStudy (anonymous):

trying it now

OpenStudy (anonymous):

ok lost

OpenStudy (mertsj):

\[2(2c)+3c=28\]

OpenStudy (mertsj):

Do you see how I replaced a with 2c?

OpenStudy (anonymous):

i got that part right

OpenStudy (mertsj):

\[4c+3c=28\]

OpenStudy (mertsj):

\[7c=28\]

OpenStudy (mertsj):

\[c=4\]

OpenStudy (mertsj):

Did you follow all that?

OpenStudy (anonymous):

ok i got it now thanks

OpenStudy (mertsj):

yw